OpenStudy (anonymous):

Help? Given cos(theta)=1/-5 and tan(theta)>0, find sin(theta)

OpenStudy (anonymous):



OpenStudy (anonymous):

find the length of the opposite side by pythagoras then take the ratio you need

OpenStudy (anonymous):

by a^2+b^2=c^2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

or by \[\sqrt{5^2-1^2}\] same thing

OpenStudy (anonymous):

4/5?

OpenStudy (anonymous):

\[\frac{ 2\sqrt{6} }{ 5 }\] ?